The square of marks age 3 years ago is equal to 6 times marks age in 9 years. What is his age now.
Answer by JoelSchwartz(130)

m=marks age currently
(m-3)^2=6(m+9)
m^2-6m+9=6m+54
m^2-12m-45=0
(m-15)(m+3)=0
(m-15)(m+3)=m^2-12m-45
m=15 or m=-3
Since marks age has to be positive his current age is 15.